OpenStudy (anonymous):

Given the polynomial 6x^3 + 4x^2 - 6x - 4, what is the value of the coefficient 'k' in the factored form? 6x^3 + 4x^2 - 6x - 4 = 2(x + k)(x - k)(3x + 2) k= ____________

OpenStudy (anonymous):

Let 2(x + k)(x - k)(3x + 2) =0 so (3x + 2) =0 i.e. x= -2/3 But (x + k)=0 or (x - k) =0 so x=-k or x=k or k=-x or k=x so k= -(-2/3) or k = -2/3 i.e. k= 2/3 or k=-2/3

OpenStudy (anonymous):

Everybody agree?

OpenStudy (anonymous):

The above answer is incorrect. 6x^3 + 4x^2 - 6x - 4 = 2(x + k)(x - k)(3x + 2) = (2x + 2k)(3x^2 + (2 - 3k)x - 2k) = 6x^3 + (4 - 6k)x^2 - 4kx + 6kx^2 + (4k - 6k^2)x - 4k^2 = 6x^3 + (4 - 6k + 6k)x^2 + (-4k + 4k - 6k^2)x - 4k^2 = 6x^3 + 4x^2 - 6k^2x - 4k^2 Comparing the constants: -4 = -4k^2 k^2 = 1 k = 1 or k -1

OpenStudy (anonymous):

so either one would be correct yea? 1 or -1? wouldn't matter?